# Intervention Efficiency and Perturbation Validation Framework: Capacity-Aware and Robust Clinical Model Selection under the Rashomon Effect

### v1

@@ 6 Conclusion
Appendix B Model Selection Stage of PVF
During the model selection stage of PVF, we simply choose the model with the highest aggregated performance from all the perturbed validation sets, as shown in Figure 5.
@@ C.1 Synthetic Dataset Experimental Setup
We vary the dataset size ($n=50$ or $100$) and the separability parameter $\mu$, which takes values uniformly spaced between 0.1 and 2.9 with a step size of 0.2, and perturbation Gaussian noise standard deviation $\sigma\in\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1}\}$ as different perturbation controls. For each hyperparameter configuration ($2\times 15\times 6=180$ combinations) in Table 1, we generate 5000 independent synthetic datasets to obtain stable comparison results and uncertainty summaries. For other hyperparameters discussed in Section 3.2 while not in Table 1, we empirically set $d=5$, $k=7$ and $M=100$ to balance the computational cost and perturbation effect. The aggregation function $\mathcal{A}$ of PVF is manually chosen to be the 25th percentile (Q1) of a model’s validation scores across perturbed validation sets, as it is well suited to high-stakes clinical contexts where worst-case generalization matters and they could correspond to the cases where all independent noises compound in a way that is most disturbing to the model’s prediction, while not being too conservative. Note that the choice of aggregation function in other applications could be decided according to specific user requirements of the attributes of the selected model though.
@@ C.2 Synthetic Dataset Experimental Results
Specifically, when the class separability is low, PVF overall selects the truly informative feature pair 42 to 65 more times than the traditional method does, across all metrics (Figure 6). This consistently positive difference confirms the superiority of PVF over the traditional method in low-separability cases, regardless of the metric. When the class separability is moderate, built on consistent positive differences, the value of difference fluctuates more, with larger values emerging at $\gamma=0.1,0.5,0.7$ (Figure 7). This could be due to a higher variability in synthetic datasets and the resulting complex relationship between IE and the classification task. F1 score/accuracy exhibit substantial positive differences, confirming that PVF is promising when these two traditional metrics are of interest in moderate-separability cases. When the class separability is high, we can witness a clearer pattern in the relationship between the performance difference and $\gamma$. That is, when $\gamma$ increases, the positive difference shrinks almost monotonically (Figure 8), with smaller $\gamma$’s (0.1, 0.3 and 0.5) enjoying a rising positive difference while larger $\gamma$’s (0.7, 0.9) suffering from a loss in positive difference. PVF with F1 Score/accuracy still maintains a large edge in the positive performance. These observations confirms that under large separability circumstances, not only is PVF consistently better than the traditional method, but also it brings large improvements when the metric of interest is IE with lower $\gamma$’s or F1 Score/accuracy.
Figure 6: Best Performance for Original Size of 50, and $\mu$ ranges from 0.1 to 0.9
Figure 7: Best Performance for Original Size of 50, and $\mu$ ranges from 1.1 to 1.9
Figure 8: Best Performance for Original Size of 50, and $\mu$ ranges from 2.1 to 2.9
@@
Although generally PVF still outperforms the traditional method for most metrics when data size is 100, a small number of cases emerge where the traditional method becomes preferable. Specifically, with moderate separability and $\gamma=0.7$ or $0.9$, the traditional method selects the true feature pair a small number of times (<<20) more than PVF (Figure 9). Nevertheless, in all other cases, PVF consistently outperforms the traditional method (Figures 10 and 11). Besides the general trend that the positive edge is larger at smaller $\gamma$’s as in data size 50, the strongest comparison occurs when the separability is high and the metrics of interest are IE with $\gamma$ (0.1, 0.3 and 0.5) or F1 Score/accuracy (Figure 11). Overall, despite a small number of slightly underperforming cases, PVF under most metrics enjoys a large edge over the traditional method with data size 100.
Figure 9: Best Performance for Original Size of 100, and $\mu$ ranges from 1.1 to 1.9
Figure 10: Best Performance for Original Size of 100, and $\mu$ ranges from 0.1 to 0.9
Figure 11: Best Performance for Original Size of 100, and $\mu$ ranges from 2.1 to 2.9

### v6

@@ 6 Acknowledgments
Appendix B Theoretical Guarantees of PVF
In this section we provide theoretical guarantees for PVF. We formalize how the PVF score behaves as the number of perturbed validation sets, the number of perturbation replicas, and the validation set size grow, and we show that PVF asymptotically selects models that are optimal for a property determined by the perturbation mechanism and the aggregation rule.
B.1 Setup
We use the notation introduced in Section 3.2. The original validation set is
$\mathcal{D}_{\mathrm{val}}=\{(x_{i},y_{i})\}_{i=1}^{n},$
where $n$ is the number of validation points, each $x_{i}$ is a feature vector, and $y_{i}$ is its label.
PVF uses a fixed perturbation mechanism that, given a feature vector $x$, produces a random perturbed feature vector $X^{\prime}$. We denote this conditional distribution by
$X^{\prime}\mid X=x\sim K(\cdot\mid x).$
In the implementation, $K$ is specified by Gaussian noise for numeric features, label-flip noise for categorical features, and decay rules for ordinal features; for the theory, we only need that $K$ is fixed and the perturbations are independent across samples and replicas.
From the original data distribution $\mathcal{P}$ on $(X,Y)$ and the perturbation rule $K$, we define the induced perturbed distribution $\tilde{\mathcal{P}}$ on $(X^{\prime},Y)$ by
$(X,Y)\sim\mathcal{P},\qquad X^{\prime}\mid X\sim K(\cdot\mid X),\qquad(X^{\prime},Y)\sim\tilde{\mathcal{P}}$.
For each perturbation index $m=1,\dots,M$, PVF constructs a perturbed validation set
$\mathcal{D}_{\mathrm{val}}^{(m)}=\bigl\{(x_{i}^{(m,j)},y_{i})\bigr\}_{i=1,\dots,n;\,j=1,\dots,k},$
where, for each fixed $i$, the replicas $\{x_{i}^{(m,j)}\}_{j=1}^{k}$ are i.i.d. drawn from $K(\cdot\mid x_{i})$, and the labels $y_{i}$ are unchanged.
For a model $f$ in a finite candidate set $F$, the evaluation metric on each perturbed set is
$P_{f}^{(m)}=\mathrm{Eval}(f,\mathcal{D}_{\mathrm{val}}^{(m)}),$
and the vector of perturbed scores is
$P_{f}=\bigl(P_{f}^{(1)},\dots,P_{f}^{(M)}\bigr)\in\mathbb{R}^{M}.$
The PVF score of model $f$ is obtained by an aggregation rule
$A_{f}=A(P_{f}),$
where $A\colon\mathbb{R}^{M}\to\mathbb{R}$ is chosen by the user (for example, $A$ can be the mean or a lower quantile). PVF then selects
$f^{\star}\in\arg\max_{f\in F}A_{f}.$
We make the following assumptions, which will be referenced explicitly in the propositions and proofs.
  1. (A1)
    $\mathcal{D}_{\mathrm{val}}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is drawn i.i.d. from a distribution $\mathcal{P}$ on $(X,Y)$.
  2. (A2)
    For each feature vector $x$, the perturbation rule is given by a fixed conditional distribution $K(\cdot\mid x)$, and all perturbations across different samples, replicas, and perturbed sets are independent.
  3. (A3)
    The set of candidate models $F$ is a finite set.
  4. (A4)
    For any fixed model $f$ and any dataset $D$, the evaluation metric $\mathrm{Eval}(f,D)$ is a bounded function of empirical statistics of $D$. In particular, for the metrics used in the paper (IE, F1, accuracy), $\mathrm{Eval}(f,D)$ depends on $D$ only through the empirical confusion matrix entries, which are empirical averages of bounded per-sample indicators. Generally, $\mathrm{Eval}(f,D)$ should be a metric that calculates a generalized mean of the performance on each data point.
  5. (A5)
    There exists a functional $\mathcal{A}$ on probability distributions on $\mathbb{R}$ such that, if $Z_{1},\dots,Z_{M}$ are i.i.d. with common distribution $L$, then
    $A(Z_{1},\dots,Z_{M})\xrightarrow[\;M\to\infty\;]{\text{a.s.}}\mathcal{A}(L).$
    This holds for standard aggregation rules including the sample mean, sample quantiles, the median, and more general continuous $L$-statistics.
We now state and prove the theoretical guarantees under these assumptions.
B.2 Convergence of PVF Scores
The first step is to understand the behavior of the PVF score $A_{f}$ for a fixed model $f$ as the sampling dimensions $M$, $k$, and $n$ grow.
For each fixed $n,k$ and realization of $\mathcal{D}_{\mathrm{val}}$, let $\mathcal{L}_{n,k}(f)$ denote the conditional distribution of $P_{f}^{(1)}=\mathrm{Eval}(f,\mathcal{D}_{\mathrm{val}}^{(1)})$ given $\mathcal{D}_{\mathrm{val}}$. Similarly, for each $n$, we will use $\mathcal{L}_{n,\infty}(f)$ for the limiting distribution of $P_{f}^{(1)}$ as $k\to\infty$.
Proposition B.1 (i.i.d. structure of perturbed scores).
Suppose (A1) and (A2) hold, and fix $\mathcal{D}_{\mathrm{val}}$, $n$, $k$, and a model $f\in F$. Then, conditional on $\mathcal{D}_{\mathrm{val}}$, the scores $P_{f}^{(1)},\dots,P_{f}^{(M)}$ are i.i.d. random variables with common distribution $\mathcal{L}_{n,k}(f)$.
Proof.
Assumption (A2) states that, for each $i$, all perturbations $\{x_{i}^{(m,j)}\}$ across different $m$ and $j$ are independent given $\mathcal{D}_{\mathrm{val}}$ and follow the same distribution $K(\cdot\mid x_{i})$. By construction, each $\mathcal{D}_{\mathrm{val}}^{(m)}$ is built from independent perturbations of $\mathcal{D}_{\mathrm{val}}$, and thus the sets $\mathcal{D}_{\mathrm{val}}^{(1)},\dots,\mathcal{D}_{\mathrm{val}}^{(M)}$ are conditionally independent and identically distributed given $\mathcal{D}_{\mathrm{val}}$. Assumption (A4) ensures that $\mathrm{Eval}(f,\cdot)$ is a deterministic and measurable function of its input dataset. Because determinism implies that $\mathrm{Eval}(f,\cdot)$ introduces no additional randomness beyond that already present in the dataset, applying $\mathrm{Eval}(f,\cdot)$ to each perturbed dataset $\mathcal{D}_{\mathrm{val}}^{(m)}$ will not create dependence between the resulting scores.
Formally, for any real intervals $I_{1},\dots,I_{k}$ and any indices $m_{1},\dots,m_{k}$, the events $\{P_{f}^{(m_{i})}\in I_{i}\}$ satisfy
$\{P_{f}^{(m_{i})}\in I_{i}\}=\{\mathcal{D}_{\mathrm{val}}^{(m_{i})}\in\mathrm{Eval}(f,\cdot)^{-1}(I_{i})\},$
where $\mathrm{Eval}(f,\cdot)^{-1}(I_{i})$ is the preimage of $I_{i}$ under the evaluation map. Since the perturbed datasets $\mathcal{D}_{\mathrm{val}}^{(1)},\dots,\mathcal{D}_{\mathrm{val}}^{(M)}$ are independent conditional on $\mathcal{D}_{\mathrm{val}}$ (by construction under Assumption (A2)), we have
$\Pr\!\left(P_{f}^{(m_{1})}\in I_{1},\dots,P_{f}^{(m_{k})}\in I_{k}\mid\mathcal{D}_{\mathrm{val}}\right)=$
$\prod_{i=1}^{k}\Pr\!\left(P_{f}^{(m_{i})}\in I_{i}\mid\mathcal{D}_{\mathrm{val}}\right).$
This shows that evaluating each $\mathcal{D}_{\mathrm{val}}^{(m)}$ through the deterministic map $\mathrm{Eval}(f,\cdot)$ preserves the i.i.d. structure established at the level of the perturbed datasets.
Therefore, the resulting scores $P_{f}^{(1)},\dots,P_{f}^{(M)}$ are conditionally i.i.d. given $\mathcal{D}_{\mathrm{val}}$ with some common distribution, which we denote by $\mathcal{L}_{n,k}(f)$. ∎
Interpretation: for a fixed validation set and perturbation parameters, PVF is aggregating many i.i.d. draws of a random performance score for each model under the chosen perturbation rule.
Proposition B.2 (Convergence in $M$).
Suppose (A1), (A2), (A4), and (A5) hold, and fix $\mathcal{D}_{\mathrm{val}}$, $n$, $k$, and $f\in F$. Then
$A_{f}=A\bigl(P_{f}^{(1)},\dots,P_{f}^{(M)}\bigr)\xrightarrow[\;M\to\infty\;]{\text{a.s.}}\mathcal{A}\bigl(\mathcal{L}_{n,k}(f)\bigr).$
Proof.
By Proposition B.1, conditional on $\mathcal{D}_{\mathrm{val}}$, the scores $P_{f}^{(1)},\dots,P_{f}^{(M)}$ are i.i.d. with common distribution $\mathcal{L}_{n,k}(f)$. Assumption (A5) states that for any i.i.d. sequence from a distribution $L$, the aggregation rule satisfies
$A(Z_{1},\dots,Z_{M})\to\mathcal{A}(L)\quad\text{almost surely as }M\to\infty.$
Applying this with $Z_{m}=P_{f}^{(m)}$ and $L=\mathcal{L}_{n,k}(f)$ yields the claimed convergence. ∎
Interpretation: when the validation set and perturbation settings are fixed, increasing the number of perturbed validation sets guarantees that PVF approximates the population-level value of the aggregation rule applied to the distribution of perturbed scores for each model.
Proposition B.3 (Convergence in $k$).
Suppose (A1), (A2), and (A4) hold, and fix $\mathcal{D}_{\mathrm{val}}$, $n$, and $f\in F$. Then, there exists a distribution $\mathcal{L}_{n,\infty}(f)$ such that
$\mathcal{L}_{n,k}(f)\xrightarrow[\;k\to\infty\;]{}\mathcal{L}_{n,\infty}(f),$
and, if (A5) also holds,
$\mathcal{A}\bigl(\mathcal{L}_{n,k}(f)\bigr)\xrightarrow[\;k\to\infty\;]{}\mathcal{A}\bigl(\mathcal{L}_{n,\infty}(f)\bigr).$
Proof.
Fix $\mathcal{D}_{\mathrm{val}}$, $n$, and $f$. By Assumption (A2), for each original sample $(x_{i},y_{i})$, the replicas $\{x_{i}^{(1,j)}\}_{j=1}^{k}$ are i.i.d. from $K(\cdot\mid x_{i})$. Assumption (A4) states that $\mathrm{Eval}(f,\mathcal{D}_{\mathrm{val}}^{(1)})$ is a bounded function of empirical statistics of $\mathcal{D}_{\mathrm{val}}^{(1)}$. By the law of large numbers, for the replicas, as $k\to\infty$, these empirical statistics of $\mathcal{D}_{\mathrm{val}}^{(1)}$ converge almost surely to their conditional expectations given $\mathcal{D}_{\mathrm{val}}$. Since $\mathrm{Eval}(f,\cdot)$ is a deterministic function of these statistics, the distribution of $P_{f}^{(1)}=\mathrm{Eval}(f,\mathcal{D}_{\mathrm{val}}^{(1)})$ also converges to a limiting distribution, which we denote by $\mathcal{L}_{n,\infty}(f)$. Assumption (A5) then implies the convergence of $\mathcal{A}(\mathcal{L}_{n,k}(f))$ to $\mathcal{A}(\mathcal{L}_{n,\infty}(f))$. ∎
Interpretation: as the number of perturbation replicas per validation point grows, the randomness arising from finite perturbation sampling disappears, and the distribution of perturbed scores for each model stabilizes to a limit.
Proposition B.4 (Convergence in $n$).
Suppose (A1), (A2), and (A4) hold. For each model $f\in F$, define
[…]
the performance value of $f$ under the perturbed population distribution $\tilde{\mathcal{P}}$. Then, as $n\to\infty$,
$\mathcal{L}_{n,\infty}(f)\xrightarrow[\;n\to\infty\;]{}\delta_{J(f)},$
where $\delta_{J(f)}$ is the probability distribution that assigns probability $1$ to the single point $J(f)$; equivalently, a random variable $Z$ has distribution $\delta_{J(f)}$ if and only if
$\Pr(Z=J(f))=1.$
If (A5) also holds, then
$\mathcal{A}\bigl(\mathcal{L}_{n,\infty}(f)\bigr)\xrightarrow[\;n\to\infty\;]{P}\mathcal{A}\bigl(\delta_{J(f)}\bigr)=:J_{A}(f).$
Proof.
Assumption (A1) implies that the empirical distribution of $\mathcal{D}_{\mathrm{val}}$ converges to $\mathcal{P}$ as $n\to\infty$. Combined with the perturbation rule in (A2), this implies that the empirical distribution of perturbed samples in each $\mathcal{D}_{\mathrm{val}}^{(1)}$ converges to $\tilde{\mathcal{P}}$. By Assumption (A4), $\mathrm{Eval}(f,\mathcal{D}_{\mathrm{val}}^{(1)})$ is a bounded function of empirical statistics of $\mathcal{D}_{\mathrm{val}}^{(1)}$. These statistics converge to their expectations under $\tilde{\mathcal{P}}$, and hence $\mathrm{Eval}(f,\mathcal{D}_{\mathrm{val}}^{(1)})$ converges in probability to $\mathrm{Eval}(f,\tilde{\mathcal{P}})=J(f)$. Therefore, the distribution $\mathcal{L}_{n,\infty}(f)$ converges to $\delta_{J(f)}$. Assumption (A5) then yields $\mathcal{A}(\mathcal{L}_{n,\infty}(f))\to\mathcal{A}(\delta_{J(f)})$. ∎
Interpretation: with a sufficiently large validation set, all randomness in the perturbed score for a fixed model collapses to a single deterministic value $J(f)$, and the limit of any reasonable aggregation of that score is just a deterministic function $J_{A}(f)$ of the same quantity.
Proposition B.5 (Unified convergence of PVF scores).
Suppose (A1)–(A5) hold. Then, for any fixed model $f\in F$,
$A_{f}\xrightarrow[\;M,k,n\to\infty\;]{P}J_{A}(f),$
[…]
Proof.
For fixed $\mathcal{D}_{\mathrm{val}}$, $n$, and $k$, Proposition B.2 implies
$A_{f}\xrightarrow[\;M\to\infty\;]{\text{a.s.}}\mathcal{A}\bigl(\mathcal{L}_{n,k}(f)\bigr).$
Proposition B.3 implies
$\mathcal{A}\bigl(\mathcal{L}_{n,k}(f)\bigr)\xrightarrow[\;k\to\infty\;]{}\mathcal{A}\bigl(\mathcal{L}_{n,\infty}(f)\bigr),$
and Proposition B.4 implies
$\mathcal{A}\bigl(\mathcal{L}_{n,\infty}(f)\bigr)\xrightarrow[\;n\to\infty\;]{P}\mathcal{A}\bigl(\delta_{J(f)}\bigr)=:J_{A}(f).$
Combining these convergences (almost sure in $M$, then in $k$, then in probability in $n$) yields the stated convergence of $A_{f}$ to $J_{A}(f)$ in probability as all three variables grow. ∎
Interpretation: for each model, as we use more perturbed validation sets, more perturbation replicas, and more validation data, the PVF score converges to a deterministic limit that depends only on the underlying data distribution, the perturbation rule, the evaluation metric, and the aggregation rule.
B.3 Consistency of PVF Model Selection
We now show that, under the same assumptions, PVF consistently selects models that maximize the objective $J_{A}(f)$.
Define
$J_{A}\colon F\to\mathbb{R},\qquad f\mapsto J_{A}(f),$
and the set of maximizers
$F_{A}^{\star}:=\{f\in F:J_{A}(f)=\max_{g\in F}J_{A}(g)\}.$
Proposition B.6 (Consistency of PVF selection).
Suppose (A1)–(A5) hold. Then the PVF-selected model $f^{\star}$ satisfies
$f^{\star}\xrightarrow[\;M,k,n\to\infty\;]{P}f_{\infty}^{\star},$
for some $f_{\infty}^{\star}\in F_{A}^{\star}$.
Proof.
By Proposition B.5, for each fixed $f\in F$,
$A_{f}\xrightarrow[\;M,k,n\to\infty\;]{P}J_{A}(f).$
Assumption (A3) states that $F$ is finite. Let $F=\{f_{1},\dots,f_{Q}\}$. For each $i\in\{1,\dots,Q\}$, define
$A_{i}:=A_{f_{i}},\qquad J_{i}:=J_{A}(f_{i}).$
Since convergence in probability holds for each coordinate individually, the whole score vector converges in probability:
$(A_{1},\dots,A_{Q})\xrightarrow[\;M,k,n\to\infty\;]{P}(J_{1},\dots,J_{Q}).$
To analyze the argmax selection rule, let
$F_{A}^{\star}:=\bigl\{i:J_{i}=\max_{j=1,\dots,Q}J_{j}\bigr\}$
be the (nonempty) set of population maximizers. For any index $i\notin F_{A}^{\star}$, we have a strict gap
$J_{i^{\star}}-J_{i}>0\qquad\text{for every }i^{\star}\in F_{A}^{\star}.$
Define
$\Delta:=\min_{i\notin F_{A}^{\star}}\,(J_{i^{\star}}-J_{i}),$
which is strictly positive.
Now, fix any $\varepsilon<\Delta/3$. We can see that, if all coordinates satisfy
$|A_{i}-J_{i}|<\varepsilon\qquad\text{for all }i=1,\dots,Q,$
then, for every $i^{\star}\in F_{A}^{\star}$ and $i\notin F_{A}^{\star}$,
$A_{i^{\star}}\geq J_{i^{\star}}-\varepsilon>J_{i}+\Delta-\varepsilon>J_{i}+2\varepsilon>A_{i}.$
Hence, whenever the coordinatewise errors are smaller than $\varepsilon$, every truly optimal model has strictly larger estimated score than every non-optimal model, and thus the maximizer of the estimated scores must lie in the true maximizer set, ensuring that we get the true optimal model:
$\arg\max_{i=1,\dots,Q}A_{i}\subseteq F_{A}^{\star}.$
Since
$\Pr\!\left(\max_{i}|A_{i}-J_{i}|>\varepsilon\right)\longrightarrow 0,$
it follows that
$\Pr\!\left(\arg\max_{i}A_{i}\subseteq F_{A}^{\star}\right)\longrightarrow 1.$
Any deterministic tie-breaking rule selects a single element of $\arg\max_{i}A_{i}$, so the selected model index converges in probability to a member of $F_{A}^{\star}$. Rewriting indices back in terms of models, we obtain:
$f^{\star}\in\arg\max_{f\in F}A_{f}\quad\Longrightarrow\quad f^{\star}\xrightarrow{P}F_{A}^{\star}.$
Thus any PVF selection rule converges in probability to a population maximizer of $J_{A}$. ∎
Interpretation: PVF is a consistent model selection procedure. In the limit of large validation sets and sufficient perturbation sampling, PVF will choose a model whose PVF score is maximal among all candidates.
B.4 PVF as Selection Toward a User-Specified Property
Finally, we interpret the limit objective $J_{A}$ as an induced property determined by the design choices in PVF.
For a fixed perturbation rule $K$, evaluation metric $\mathrm{Eval}$, and aggregation rule $A$, we define the induced property
[…]
This property assigns to each model $f$ a scalar that reflects its performance under the perturbed distribution $\tilde{\mathcal{P}}$, summarized through the aggregation rule $A$.
Proposition B.7 (PVF selects toward the induced property).
Suppose (A1)–(A5) hold. Let $\Phi_{A,K}(f):=J_{A}(f)$ be the induced property described above. Then
$f^{\star}\xrightarrow[\;M,k,n\to\infty\;]{P}f_{\mathrm{opt}},$
for some
$f_{\mathrm{opt}}\in\arg\max_{f\in F}\Phi_{A,K}(f).$
Proof.
This is a direct restatement of Proposition B.6, since $\Phi_{A,K}(f)$ is defined to be $J_{A}(f)$ for each $f\in F$. ∎
Interpretation: once a user fixes a perturbation mechanism and an aggregation rule, these choices implicitly define a property $\Phi_{A,K}$ of interest (for example, robustness to specific perturbations, emphasis on lower-tail performance, or any other stability notion encoded by $K$ and $A$). The propositions in this section show that, under the stated assumptions, PVF is guaranteed to select models that maximize this induced property.
@@ C.1 Synthetic Dataset Experimental Setup
We vary the dataset size ($n=50$ or $100$) and the separability parameter $\mu$, which takes values uniformly spaced between 0.1 and 2.9 with a step size of 0.2, and perturbation Gaussian noise standard deviation $\sigma\in\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1}\}$ as different perturbation controls. For each hyperparameter configuration ($2\times 15\times 6=180$ combinations) in Table 1, we generate 5000 independent synthetic datasets to obtain stable comparison results and uncertainty summaries. For other hyperparameters discussed in Section 3.2 while not in Table 1, we empirically set $d=5$, $k=7$ and $M=100$ to balance the computational cost and perturbation effect. The aggregation function $\mathcal{A}$ of PVF is manually chosen to be the 25th percentile (Q1) of a model’s validation scores across perturbed validation sets for demonstration, as it is well suited to high-stakes clinical contexts where worst-case generalization matters and they could correspond to the cases where all independent noises compound in a way that is most disturbing to the model’s prediction, while not being too conservative. Note that the choice of aggregation function in other applications could be decided according to specific user requirements of the attributes of the selected model, as Proposition B.7 has stated that ideally PVF should be able to select the optimal model toward any user-defined property.
@@ C.2 Synthetic Dataset Experimental Results
Specifically, when the class separability is low, PVF overall selects the truly informative feature pair 42 to 65 more times than the traditional method does, across all metrics (Figure 6). This consistently positive difference confirms the superiority of PVF over the traditional method in low-separability cases, regardless of the metric.
Figure 6: The best performance of PVF on synthetic datasets ($n=50,\mu\in[0.1,0.9]$) under different metrics, when $\sigma$ is tuned. PVF identifies the truly informative feature pair more frequently across all evaluation metrics.
When the class separability is moderate, built on consistent positive differences, the value of difference fluctuates more, with larger values emerging at $\gamma=0.1,0.5,0.7$ (Figure 7). This could be due to a higher variability in synthetic datasets and the resulting complex relationship between IE and the classification task. F1 score/accuracy exhibit substantial positive differences, confirming that PVF is promising when these two traditional metrics are of interest in moderate-separability cases.
When the class separability is high, we can witness a clearer pattern in the relationship between the performance difference and $\gamma$. That is, when $\gamma$ increases, the positive difference shrinks almost monotonically (Figure 8), with smaller $\gamma$’s (0.1, 0.3 and 0.5) enjoying a rising positive difference while larger $\gamma$’s (0.7, 0.9) suffering from a loss in positive difference. PVF with F1 Score/accuracy still maintains a large edge in the positive performance. These observations confirms that under large separability circumstances, not only is PVF consistently better than the traditional method, but also it brings large improvements when the metric of interest is IE with lower $\gamma$’s or F1 Score/accuracy.
@@
Although generally PVF still outperforms the traditional method for most metrics when data size is 100, a small number of cases emerge where the traditional method becomes preferable. Specifically, with moderate separability and $\gamma=0.7$ or $0.9$, the traditional method selects the true feature pair a small number of times (<<20) more than PVF (Figure 9).
Nevertheless, in all other cases, PVF consistently outperforms the traditional method (Figures 10 and 11). Besides the general trend that the positive edge is larger at smaller $\gamma$’s as in data size 50, the strongest comparison occurs when the separability is high and the metrics of interest are IE with $\gamma$ (0.1, 0.3 and 0.5) or F1 Score/accuracy (Figure 11). Overall, despite a small number of slightly underperforming cases, PVF under most metrics enjoys a large edge over the traditional method with data size 100.
Figure 7: The best performance of PVF on synthetic datasets ($n=50,\mu\in[1.1,1.9]$) under different metrics, when $\sigma$ is tuned. PVF continues to outperform the traditional method across all metrics, with notably higher improvements when $\gamma\in\{0.1,0.5,0.7\}$. Variability increases relative to low-separability conditions, reflecting more complex interactions between IE and classification difficulty.
Figure 8: The best performance of PVF on synthetic datasets ($n=50,\mu\in[2.1,2.9]$) under different metrics, when $\sigma$ is tuned. PVF still outperforms the traditional method, but the magnitude of improvement declines almost monotonically as $\gamma$ increases.
Figure 9: The best performance of PVF on synthetic datasets ($n=100,\mu\in[1.1,1.9]$) under different metrics, when $\sigma$ is tuned. Although PVF generally outperforms the traditional method, a few cases where the traditional method selects the true feature pair slightly more often appear when $\gamma\in\{0.7,0.9\}$
Figure 10: The best performance of PVF on synthetic datasets ($n=100,\mu\in[0.1,0.9]$) under different metrics, when $\sigma$ is tuned. PVF consistently selects the informative feature pair more frequently across all metrics, with larger advantages at smaller $\gamma$ values.
Figure 11: The best performance of PVF on synthetic datasets ($n=100,\mu\in[2.1,2.9]$) under different metrics, when $\sigma$ is tuned. PVF exhibits its strongest improvements under high separability, especially when $\gamma\in\{0.1,0.3,0.5\}$ and for F1 Score/accuracy.